OpenStudy (anonymous):

Subtract and simplify:2 over 3x+5 minus 1 over 6x+10

OpenStudy (pfenn1):

\[\frac{2}{3x+5}-\frac{1}{6x+10}=\frac{2}{3x+5}-\frac{1}{2(3x+5)}\]Can you solve it from here. (Hint: look for LCD)

OpenStudy (anonymous):

3x+5 would be the common denominator ?

OpenStudy (pfenn1):

You would also need the 2. The common denominator=2(3x+5). The least common denominator of two or more non-zero denominators is actually the smallest whole number that is divisible by each of the denominators. There is a lot of good discussion on how to find them here : http://www.helpwithfractions.com/least-common-denominator.html

OpenStudy (pfenn1):

so the expression would now be\[\frac{2(2)-1}{2(3x+5)}\]Can you simplify it further?